# Quantum Computational Insurance and Actuarial Science

### v1

@@ Abstract
In recent years, quantum computation has been rapidly advancing, driving a technological revolution with significant potential across various sectors, particularly in finance. Despite this, the insurance industry, an essential tool for mitigating unforeseen risks and losses, has received limited attention. This paper provides an initial exploration into the realm of quantum computational insurance and actuarial science. After introducing key insurance models and challenges, we examine quantum algorithms designed to address complex insurance issues. Our study includes experimental and numerical demonstrations of quantum applications in non-life insurance, life insurance, and reinsurance. Additionally, we explore the timeline for quantum insurance, the development of quantum-enhanced insurance products, and the challenges posed by quantum computational advancements.
I Introduction
@@
Insurance serves as a financial mechanism that helps individuals avoid potential losses and hedge against risks from specific sources [1]. It encompasses three main branches: non-life insurance, life insurance, and risk management. In the branch of non-life insurance, the central tasks are to price the premiums paid by an individual policyholder, to model the claims for an insurer to afford when an insured event happens, and to estimate the liability considering the Incurred But Not Reported (IBNR) and Reported But Not Paid (RBNP) claims [2]. While the life insurance branch pays attention to insure life against disability-and-death risk by mortality forecasting [3, 4, 5], as well as against live-too-long risk by life annuity pricing [3, 6]. In the risk management branch, reinsurance and risk indicators are common tools for transferring and quantifying the risks [7]. Ultimately, the core tasks of insurance practice and actuarial science revolve around modeling and evaluating random risky events [8]. To achieve these objectives, a variety of computational methods are employed.
Nevertheless, the traditional insurance industry and academia are encountering challenges due to the rapidly growing size of data and increasingly complex models, which will cost exponentially increased computational resources. One exciting answer to these problems is quantum computation [9], which has been widely applied in chemistry [10], biology [11, 12], materials [13], drug design [14], and especially finance [15, 16, 17]. Specifically, the quantum simulation of continuous-time stochastic processes is introduced to model the collective risk and the insurance cash flow [18, 19]. However, quantum computational insurance is still in its infancy compared with other financial branches: few quantum algorithms focus on insurance practice, and no systematic study on actuarial science has been proposed [15, 16, 17].
@@
This article provides an early yet insightful exploration of addressing challenges in insurance practice and actuarial science with quantum computation, We respectively review the main problems in non-life insurance, life insurance, and risk management. Then quantum algorithms for these tasks are discussed. Moreover, we also discuss them with respect to the noisy intermediate-scale quantum (NISQ) era [20] and fault-tolerant quantum computation (FTQC) era [21] based on their requirement for quantum resources, offering a timetable for quantum computational insurance.
@@
We also numerically and experimentally study several quantum insurance applications corresponding to mainstream insurance branches, providing further evidence of the feasibility of quantum computational methods for realistic insurance problems. Initially, we present numerical simulations for the excess evaluation problem. Following this, we experimentally demonstrate solutions for reinsurance type allocation and the Lee-Carter Model [4] using the superconducting quantum processor “Wukong” [22]. This marks the first implementation of an insurance-specific quantum algorithm on a quantum processor using real datasets.
@@
Beyond the applicability and demonstrability of quantum insurance, we further explore the influence and impact of quantum technology on the insurance industry. We consider the three main developmental stages of quantum insurance and highlight how quantum computation is driving the creation of new types of insurance products. Despite these advancements, some fundamental challenges still need to be addressed.
II Basics in quantum computation
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (1)
An $n$-qubit quantum logic gate can process the $2n$-dimensional information in one step, outperforming classical logic gates. It can also cause entanglements or interference between qubits. These features make quantum computation potential to achieve speedups over classical computations. Quantum algorithms based on these principles have been designed for various regions including chemistry, biology, finance, etc.
III Problems related to insurance
III.1 Non-life Insurance
@@ III.3 Risk Management
IV Quantum algorithms for insurance
As aforementioned, modeling stochastic events is the core task of insurance and actuarial science. Quantum stochastic modeling involves embedding classical stochastic information into a quantum computer, processing the embedded data, and extracting the associated classical information for practical utilization. Efficiency in these steps is required for quantum advantages. Fortunately, quantum stochastic modeling could reduce the space exponentially with a further quadratic speedup. Firstly, statistical distributions, as well as initial states of stochastic processes, can be prepared in a quantum superposition, wherein the intrinsic symmetry of the underlying distribution can be employed to reduce the time complexity and mitigate the classical quantum input bottleneck. Secondly, for a stochastic process, the Markov dynamics described by a transition matrix can be simulated with quantum operations, herein the memory-less property can be employed to extend to continuous-time processes at a low cost. While the non-Markovian behaviors are related to an open quantum system, whose simulation has been widely studied [42, 43]. Finally, to evaluate the desired information from the expectation of moments, we implement the quantum-enhanced Monte-Carlo integration (QMC). It applies amplitude estimation to the quantum samplings described above with a quadratic accelerate [44, 45].
@@ IV Quantum algorithms for insurance
While in the FTQC era, where quantum circuits and the qubit number are greatly improved, it is possible to implement quantum machine learning algorithms based on linear algebra. The core subroutine in these algorithms is the solution of linear equations using the widely applied HHL algorithm [47, 48]. For instance, using quantum computers, The vision transformer [49], which is successfully applied on natural language processing tasks, can achieve speedups [50]. the singular value decomposition to extracting features can also be accelerated with quantum algorithms [51].
@@
V Results
In this part, we numerically and experimentally study several insurance problems. First, we develop a quantum algorithm to solve the excess evaluation problem. The quantum computing framework QPanda [56] is used for this task. Then we solve the reinsurance type allocation problem and the Lee-Carter model on “Wukong” [22], which is a high-performance superconducting quantum computer with the average relaxation time $T_{1}=14.84\mu s$ and $T_{2}=1.85\mu s$. The average fidelity of single-qubit gates and two-qubit gates are respectively $99.7\%$ and $96.3\%$.
V.1 Quantum excess evaluation
This part shows the numerical simulation results of the policy excess introduced before. We aim to model the payment for the reinsurer to pay $\mathbb{E}[\max(0,X-M)]$, or the discrete form:
$\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}\max(0,X_{j}-M(X_{j})),$ (12)
where $N_{c}$ is the total claims. We assume $X_{j}$ to be the $j$-th loss followed the distribution
$f(x,\mu,\sigma)=\frac{1}{x\sigma\sqrt{2\pi}}e^{-\frac{(\ln x-\mu)^{2}}{2\sigma^{2}}}.$ (13)
$M(X_{j})$ is the payment, which is a non-decreasing function:
@@ V.1 Quantum excess evaluation
As the policy number grows and the underlying distribution varies, a theoretical evaluation turns to be intractable and a numerical simulation demands a huge computation source consumption [2].
We developed a quantum algorithm for this problem, which is detailed in Sec. VII.1. In our simulation, we set $\mu=0$ and $\sigma=1$. Due to the convergence properties of the integral, it can be truncated with a certain level of precision. In this simulation, the excess $X_{j}$ is varied from 0 to 10 and The simulation results are shown in Fig. 1. As the number of qubits used to represent the distribution increases, the evaluated expectation becomes more accurate and convergent.
[…]
Figure 1: Simulation results of excess evaluation. The green and yellow lines represent the theoretical payment and its truncated version, respectively, demonstrating that truncating to $X\in[0,10]$ provides adequate accuracy. The blue dashed line shows the simulated results as a function of the number of qubits used. As the number of qubits increases, the simulation results converge towards greater accuracy.
V.2 Quantum reinsurance type allocation
@@
Herein the risk $R_{j}$ can be either a vector of historical data or a random variable from stochastic simulation, and the proportion $p$ satisfying $10<p<$ 1 represents the average allocation probability and is assumed to be fixed (for example, $p=0.5$ as in [60]).
We solve the case that the data is generated randomly. Noting that the loss function can be converted into an en expectation value of one Ising model through the standard transformation $x_{j}\rightarrow\frac{I-Z_{j}}{2}$ as
$H_{cost}=\sum_{j,k=1}^{n}v_{jk}\frac{I-Z_{j}}{2}\frac{I-Z_{k}}{2},$ (16)
[…]
We applied 6 superconducting qubits, illustrated in Fig. 2(a), for this task. To satisfy the constraint, instead of a traditional penalty function method, we utilize a more efficient problem-specified ansatz consisting of two parts: a parameterized initialization layer to prepare a quantum state satisfying the constraint, and repeating layers of reconfigurable Beam splitter (RBS) gates to transform the state while keeping the constraint unviolated [61, 62]. The RBS gate is $\text{diag}\left\{1,\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix},1\right\}$, which interacts between $|01\rangle$ and $|10\rangle$ while keeps $|00\rangle$ and $|11\rangle$ unchanged. A quantum circuit implementation of this gate is shown in Fig. 2(b). The number of RBS layers is set to be 3 and the whole ansatz after some simplifications is shown in Fig. 2(c).
[…]
(a)(b)(c)
Figure 2: (a) The 6 superconducting qubits used in the quantum computer “Wukong”. (b) The quantum circuit for the RBS gates. (c) The quantum circuit for the ansatz, with 6 qubits and 3 RBS layers, after some simplifications.
We apply a measurement-based error mitigation strategy to combat quantum hardware noise. While the theoretical output maintains the constraints, any deviation due to hardware noise is corrected by excluding erroneous shots from measurements.
@@
V.3 Quantum Lee-Carter solver
@@
Here, we report experimental results for solving the Lee-Carter model described in Eq. (7). Given a list of data $m_{x,t}$, the core task is to estimate $\beta_{x}$ and $\kappa_{t}$. This is usually done by first constructing the matrix $D$ whose elements are $D_{x,t}=\ln m_{x,t}-\alpha_{x}$. Then a singular value decomposition is performed on $D$, and the left(right) singular vector related to the maximum singular value is associated to $\beta_{x}(\kappa_{t})$. Then these data can be used for further predictions. We apply the quantum singular value decomposition algorithm [63]. A brief introduction of this algorithm is shown in Sec. VII.2.
@@
We apply real-world data obtained from  [^1] https://www.mortality.org. The year $t$ is from $2014$ to $2017$ and the age $x$ is from 00 to $10-14$. Then the matrix size of $D$ is set as $4\times 4$. A total of 4 superconducting qubits are used and a total of 200 optimization steps are evaluated. The experimental results are shown in Fig. 4. It shows that the loss function decreases with the optimization steps. During this process, both the Frobenius norm distance and the Kullback-Leibler (K-L) divergence between the theoretical singular vector and our estimated singular vector also decrease, indicating the successful performance of the target singular value decomposition.
@@
Figure 4: Experimental results of the Lee-Carter model. (a), (b), (c), and (d) respectively represent the evolution of the loss function, the Frobenius norm distance, and the K-L divergence (left and right) between the theoretical singular vector and our optimized singular vector over the epochs.
VI Conclusion and discussion

### v2

@@ Abstract
In recent years, quantum computation has been rapidly advancing, driving a technological revolution with significant potential across various sectors, particularly in finance. Despite this, the insurance industry, an essential tool for mitigating unforeseen risks and losses, has received limited attention. This paper provides an initial exploration into the realm of quantum computational insurance and actuarial science. After introducing key insurance models and challenges, we discuss quantum algorithms that can address insurance problems based on their mathematical nature. Our study includes experimental and numerical demonstrations of quantum applications in non-life insurance, life insurance, and reinsurance. Additionally, we explore the timeline for quantum insurance, the development of quantum-enhanced insurance products, and the challenges posed by quantum computational advancements. This work systematically constructs the connection between quantum computation and the insurance industry, enhancing the development of insurance while promoting the application of quantum computation to more realistic problems.
keywords:
quantum computation, insurance, quantum machine learning, superconducting quantum compuiters
1 Introduction
@@
Insurance is the financial means helping people to avoid potential loss and to hedge against risk from particular sources [1]. What lies in the core and crux of insurance practice and actuarial science is to model and evaluate random risky events [2]. In the branch of non-life insurance, The central tasks are to price the premiums paid by an individual policyholder, to model the claims for an insurer to afford when an insured event happens, and to estimate the liability considering the Incurred But Not Reported (IBNR) and Reported But Not Paid (RBNP) claims [3]. Therein, Monte-Carlo simulation, regression, and neural network are introduced to solve complicated statistics mathematics and stochastic process problems [4, 5, 6]. In contrast, the life insurance branch pays attention to insure life against disability-and-death risk by mortality forecasting [7, 8, 9], as well as against live-too-long risk by life annuity pricing [7, 10]. This can be solved by mortality forecasting models, financial/mortality stochastic processes simulations, and multi-period portfolio optimization [11, 12, 13]. As for the risk management branch, reinsurance and risk indicators are common tools to transfer and quantify the risks, respectively [14]. Consequently, stochastic process catching large claims and risk portfolio optimization are both emphasized when studying the reinsurance problem [7, 11, 15, 16]. Nevertheless, the traditional insurance industry and academia are encountering challenges due to the rapidly growing data size and increasingly complex models, which will require exponentially increased computational resources [11].
One exciting answer to these problems is quantum computation [17], which has been widely applied in chemistry [18], biology [19, 20], materials [21], drug design [22], and especially finance [23, 24, 25]. Specifically, applying quantum computing methods to insurance problems has also been investigated, including the insurance cash flow modeling [26], collective risk model [27], insurance capital modeling [28, 29], risk identification [30, 31], insurance claim fraud determination [32], etc. However, quantum computational insurance is still in its infancy compared with other financial branches: Firstly, existing works mainly focus on applying quantum computation to specific insurance problems while a systematic study on the general framework has not been established. Secondly, experimental demonstrations have not been implemented on current quantum computers to pave the way for the applicability and practicality of quantum insurance. Thirdly, while empirical analysis of these models on real-world data sets is essential for any insurance problems, relevant research is still absent.
We argue that the above phenomena and challenges originate from the significant difference between insurance mathematics and many other financial problems. From the theoretical aspect, as statistical and actuarial models have been developed and applied to manage risk for hundreds of years, critical analysis and theoretical interpretability have been proven fundamental to understanding their stochastic basis [33]. The fancy relationships between quantum information and the stochastic nature of risk theory, including multivariate statistical analysis, risk processes, and risks in the collective [34], need to be released. From the practical aspect, realistic data-based empirical analysis is essential to calibrate and validate insurance and risk models [35, 36, 37], while efficient data encoding methods are still challenging for quantum computing, especially in its current Noisy Intermediate Scale Quantum (NISQ) era [38, 39, 40, 41].
@@
This article provides an early yet insightful exploration of addressing challenges in insurance practice and actuarial science with quantum computation. To provide an intuitive perception of insurance, we first review the problems in the main branches of non-life insurance, life insurance, and risk management. Then we categorize the problems as stochastic modeling tasks, optimization tasks, and machine learning-based tasks, and quantum algorithms to solve these tasks with potential speedup are introduced. Moreover, we also discuss them with respect to the NISQ era [42] and fault-tolerant quantum computation (FTQC) era [43] based on their requirement for quantum resources, offering a timetable for quantum computational insurance.
@@
We also numerically and experimentally study several quantum insurance applications corresponding to mainstream insurance branches. Initially, for the policy excess problem, we designed the quantum excess evaluation algorithm to model the (stochastic) payment and performed numerical simulations. Following this, we experimentally demonstrate solutions for the reinsurance type allocation problems with variational optimization algorithms and the Lee-Carter model [8] with machine learning-based algorithms on the superconducting quantum processor “Wukong” [44]. Notably, solving the Lee-Carter model for mortality forecasting is the first implementation of an insurance-specific quantum algorithm on a quantum processor using real datasets. These demonstrations and empirical analyses provide further evidence of the feasibility of quantum computational methods for realistic insurance problems.
@@
Beyond the applicability and demonstrability of quantum insurance, we explore its influence and impact on the insurance industry. We consider the three main developmental stages of quantum insurance and highlight how quantum computation is driving the creation of new types of insurance products. Despite these advancements, some fundamental challenges still need to be addressed.
Combining these aspects, our work introduces the connection between quantum computation and the insurance industry. This can pave the way for the development of the insurance industry with quantum computation while promoting the application of quantum computation to more realistic problems.
2 Problems related to insurance
2.1 Non-life Insurance
@@ 2.3 Risk Management
3 General framework for quantum computational insurance
In this section, we establish a general framework for quantum computational insurance. After a brief introduction to the basics of quantum computation, we explore the connections between insurance problems and quantum computation. Firstly, the fundamental concept of probability distributions can be efficiently encoded on quantum states by state preparation algorithms. Secondly, the essential model of risk processes can be simulated and analyzed by quantum stochastic simulation algorithms. Thirdly, the more complicated collective and portfolio of risk processes can be optimized by quantum optimization algorithms. Finally, the rapidly growing data-driven insurance models can be trained by quantum machine learning models. Following them, more specific algorithm implementations are left to the next section.
3.1 Basics in quantum computation
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (9)
An $n$-qubit quantum logic gate can process the $2^{n}$-dimensional information in one step, outperforming classical logic gates. It can also cause entanglements or interference between qubits. These features make quantum computation potential to achieving speedups over classical computations. Quantum algorithms based on these principles have been designed for various regions, including chemistry, biology, finance, etc.
3.2 Quantum algorithms for insurance
This part discusses quantum algorithms for insurance problems. An overview is shown in Fig. 1.
Figure 1: Quantum algorithms for insurance problems, which are categorized as stochastic modeling tasks, machine learning tasks, and optimization tasks. The policy excess problem, the Lee-Carter model, and the reinsurance type allocation problem are typical examples in these classes and are demonstrated numerically and experimentally in the next section.
As aforementioned, modeling stochastic events is the core task of insurance and actuarial science. For instance, CRM and the policy excess problem introduced before. Quantum stochastic modeling involves embedding classical stochastic information into a quantum computer, simulating the stochastic process, and extracting the associated classical information for practical utilization. Efficiency in these steps is required for quantum advantages. Fortunately, quantum stochastic modeling could reduce the space exponentially with a further quadratic speedup. Firstly, statistical distributions, as well as initial states of stochastic processes, can be prepared in a quantum superposition, wherein the intrinsic symmetry of the underlying distribution can be employed to reduce the time complexity and mitigate the classical quantum input bottleneck. Secondly, for a stochastic process, the Markov dynamics described by a transition matrix can be simulated with quantum operations, herein the memory-less property can be employed to extend to continuous-time processes at a low cost. While the non-Markovian behaviors are related to an open quantum system, whose simulation has been widely studied [63, 64]. Finally, to evaluate the desired information from the expectation of moments, we implement the quantum-enhanced Monte Carlo integration (QMC). It applies amplitude estimation to the quantum samplings described above with a quadratic accelerate [65, 66].
@@ 3.2 Quantum algorithms for insurance
While in the FTQC era, where quantum circuits and the qubit number are greatly improved, it is possible to implement quantum machine learning algorithms based on linear algebra. The core subroutine in these algorithms is the solution of linear equations using the widely applied HHL algorithm [68, 69]. For instance, using quantum computers, The vision transformer [70], which is successfully applied on natural language processing tasks, can achieve speedups [71]. The singular value decomposition (SVD) to extracting features can also be accelerated with quantum algorithms [72].
@@
4 Results
In this part, we numerically and experimentally study insurance problems covering the main branches as non-life insurance, life insurance, and reinsurance. Based on the discussion of quantum algorithms, we respectively solve them with quantum stochastic modeling algorithms, quantum optimization algorithms, and quantum machine learning-based algorithms. First, we develop a quantum algorithm to model the (stochastic) payment in the excess evaluation problem. The quantum computing framework QPanda [77] is used for this task. Then we solve the reinsurance type allocation problem with variational quantum optimization algorithms and the Lee-Carter model with quantum singular value decomposition algorithms on “Wukong” [44], which is a high-performance superconducting quantum computer with the average relaxation time $T_{1}=14.84\mu s$ and $T_{2}=1.85\mu s$. The average fidelity of single-qubit gates and two-qubit gates are respectively $99.7\%$ and $96.3\%$.
4.1 Quantum excess evaluation
In this part, we numerically solve the policy excess problem, which aims to model the payment for the reinsurer to pay $\mathbb{E}[\max(0,X-M)]$. In this task, we consider the loss following the distribution
$f(x,\mu,\sigma)=\frac{1}{x\sigma\sqrt{2\pi}}e^{-\frac{(\ln x-\mu)^{2}}{2\sigma^{2}}},$ (12)
where $\mu=0$ and $\sigma=1$ are used. The threshold $M(x)$ is set to be a non-decreasing function
@@ 4.1 Quantum excess evaluation
In this case, the payment can be re-expressed as $R_{\text{theory}}=\int_{1}^{\infty}f(x)(x-M(x))dx$. Due to the convergence properties of the integral, it can be truncated with a certain level of precision. In this simulation, the excess $x$ is varied from 0 to 10 and the corresponding truncated payment is $R_{\text{truncate}}=\int_{x=1}^{10}g(x)(x-M(x))dx$, where $g(x)=\frac{f(x)}{\int_{x=0}^{10}f(x)dx}$. In real simulations, we also need to discrete the range into $N=2^{n}$ points (claims), The loss of each claim is $X_{j},0\leq j\leq N-1$. Then the payment becomes $R_{\text{discrete}}=\sum_{X_{j}\geq 1}f(X_{j})(X_{j}-M(X_{j}))$.
We develop a quantum algorithm for this problem, which uses $n+2$ qubits according to the discrete of the excess range. The quantum algorithm starts with preparing the quantum state representing the distribution of loss, then a quantum subtracter is applied to evaluate each $X_{j}-M(X_{j})$ in parallel. Then a sequence of controlled rotations is applied to establish the connection between the payment and the results when measuring auxiliary qubits. Details of the algorithm can be found in Appendix A.
The numerical simulation result of this task is shown in Fig. 2. Several conclusions can be observed. Firstly, it works when we truncate for large $x$. Secondly, the correctness of the algorithm grows with the number of qubits. Thirdly, the algorithm can characterize the excess evaluation problem well.
Figure 2: Simulation results of excess evaluation. $R_{\text{theory}}$ and $R_{\text{truncate}}$ represent the theoretical and the truncated one, demonstrating that only considering $x\in[0,10]$ provides adequate accuracy. $R_{\text{discrete}}$ is the payment evaluated with our algorithm, showing that as the number of qubits increases, the accuracy is improved and converged to $R_{\text{truncate}}$
4.2 Quantum reinsurance type allocation
@@
Herein the risk $R_{j}$ can be either a vector of historical data or a random variable from stochastic simulation, and the proportion $p$ satisfying $0<p<1$ represents the average allocation probability and is assumed to be fixed (for example, $p=0.5$ as in [81]).
Expand the above variance, we have:
$\text{Var}[R_{\text{total}}]=\sum_{j}\text{Var}[x_{j}R_{j}]+\sum_{j\neq k}\text{Cov}[x_{j}R_{j},x_{k}R_{k}],$ (15)
where $\text{Cov}[X,Y]$ is the covariance between $X$ and $Y$. Using $\text{Var}[X]=\text{Cov}[X,X]$, the equation becomes:
$\text{Var}[R_{\text{total}}]=\sum_{jk}\text{Cov}[x_{j}R_{j},x_{k}R_{k}]=\sum_{jk}x_{j}x_{k}\text{Cov}[R_{j},R_{k}].$ (16)
Denote the covariance matrix $V$ with elements $v_{jk}=\text{Cov}[R_{j},R_{k}]$, We have:
$\text{Var}[R_{\text{total}}]=\sum_{jk}x_{j}v_{jk}x_{k}=x^{T}Vx.$ (17)
In this task, we randomly generate the covariance matrix, and a VQA is applied. We generate the ansatz with some parameterized unitary $|x(\bm{\theta})\rangle=U(\bm{\theta})|0\rangle$. To evaluate the variance with quantum measurements, we apply the map $x_{j}\to(I-Z_{j})/2$ such that $x_{j}=\langle x_{j}|(I-Z_{j})/2|x_{j}\rangle,x_{j}=0,1$. Then the variance becomes:
$\text{Var}[R_{\text{total}}]=\langle x|H_{\text{cost}}|x\rangle,\qquad H_{\text{cost}}=\sum_{j,k=1}^{n}v_{jk}\frac{I-Z_{j}}{2}\frac{I-Z_{k}}{2},$ (18)
Then it is seen that evaluating the variance becomes a measure of the expectation value of the Hamiltonian $H_{\text{cost}}$. Finding $x$ satisfying $\sum_{j}x_{j}=pn$ that minimizes the variance becomes finding the eigenstate corresponding to the smallest eigenvalue of the Hamiltonian’s specific subspace. Finally, we can apply a classical optimizer to optimize the results.
We applied 6 superconducting qubits, illustrated in Fig. 3(a), for this task. To satisfy the constraint, instead of a traditional penalty function method, we utilize a more efficient problem-specified ansatz consisting of two parts: a parameterized initialization layer to prepare a quantum state satisfying the constraint, and repeating layers of reconfigurable Beam splitter (RBS) gates to transform the state while keeping the constraint unviolated [82, 83, 84]. The RBS gate is $\text{diag}\left\{1,\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix},1\right\}$, which interacts between $|01\rangle$ and $|10\rangle$ while keeps $|00\rangle$ and $|11\rangle$ unchanged. A quantum circuit implementation of this gate is shown in Fig. 3(b). The number of RBS layers is set to be 3 and the whole ansatz after some simplifications is shown in Fig. 3(c).
Figure 3: Setup for the reinsurance type allocation problem. (a) The 6 superconducting qubits used in the quantum computer “Wukong”. (b) The quantum circuit for the RBS gates. (c) The quantum circuit for the ansatz, with 6 qubits and 3 RBS layers, after some simplifications.
While the theoretical output maintains the constraints, any deviation due to hardware noise is corrected by excluding erroneous shots from measurements. We apply a measurement-based error mitigation strategy to combat quantum hardware noise. It first runs several quantum circuits to benchmark the probability of flips when measuring 0 and 1. Then after the measurement outcome is obtained, we perform a reverse operation to get the noise-mitigated result.
@@
4.3 Quantum Lee-Carter solver
@@
Here, we report experimental results for solving the Lee-Carter model described in Eq. (4). Given a list of data $m_{x,t}$, the core task is to estimate $\beta_{x}$ and $\kappa_{t}$. This is usually done by first constructing the matrix $D$ whose elements are $D_{x,t}=\ln m_{x,t}-\alpha_{x}$. Then a singular value decomposition is performed on $D$, and the left(right) singular vector related to the maximum singular value is associated with $\beta_{x}(\kappa_{t})$. Then these data can be used for further predictions. We apply real-world data obtained from  [^1] https://www.mortality.org. The year $t$ is from $2014$ to $2017$ and the age $x$ is from 0 to 101410-14. Then the matrix size of $D$ is set as $4\times 4$.
We apply the quantum singular value decomposition algorithm [72]. A brief introduction to this algorithm is shown in Appendix B. A total of 4 superconducting qubits are used, and a total of 200 optimization steps are evaluated. Similar to the previous experiment, we directly optimized the parameters on a quantum computer and used those parameters for benchmarks. Denote the optimized left and right vectors in the optimization process as $\beta$ and $\kappa$, then to characterize the optimization process, we observe the Frobenius norm distance between the data matrix $D$ and our constructed matrix $\beta\kappa$, $\|D-\beta\kappa\|_{F}$, and the Kullback-Leibler (K-L) divergence between the optimized vector and the target one, $D_{\text{KL}}(\beta\|\beta_{x})$ and $D_{\text{KL}}(\kappa\|\kappa_{t})$, where $D_{\text{KL}}(p\|q)=\sum_{i}p_{i}\log(p_{i}/q_{i})$.
@@
The experimental results are shown in Fig. 5. It shows that the loss function decreases with the optimization steps. During this process, both the Frobenius norm distance between the data matrix and our constructed one and the K-L divergence between the theoretical singular vector and our estimated singular vector also decrease, indicating the successful performance of the target singular value decomposition. Note that this is the first time an insurance-specific real-world data problem is experimentally solved with real quantum computers.
@@
Figure 5: Experimental results of the Lee-Carter model. (a) Evolution of the loss function in the optimization process. (b) Evolution of the Frobenius norm distance between the data matrix $D$ and our constructed matrix $\beta\kappa$, where $\beta$ and $\kappa$ are the left and right vectors in the optimization process. (c) Evolution of the K-L divergence between the theoretical vector $\beta_{x}$ and $\beta$. (d) Evolution of the K-L divergence between the theoretical vector $\kappa_{t}$ and $\kappa$.
5 Conclusion and discussion